PHOTODISSOCIATION SPECTRA OF CYCLIC AND NON-CYCLIC NUCLEOTIDES

\underline{SYDNEY H. KAUFMAN}, JESSE C. MARCUM, and J. MATHIAS WEBER, JILA, NIST, and Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309.

We have studied the UV photodepletion and photofragment action spectra of deprotonated adenosine monophosphate and guanosine monophosphate as well as their cyclic analogs. We observe the same anionic fragments as in earlier experiments on collision-induced dissociation of deoxyribonucleotides, although their relative intensities are quite different, especially with respect to the abundance of the deprotonated base anions. The suppression and amplification of spectral features provides information about the threshold energies for the active fragment channels.